OpenStudy (anonymous):

Which expression describes the rule of this sequence? 7, 14, 21, 28, 35…

OpenStudy (anonymous):

@austinL

OpenStudy (austinl):

7a = 7,14,21,28,35,... Is this what you mean?

OpenStudy (anonymous):

You're looking for a pattern here. Start by seeing if each value is just adding the same value to the previous one. You can do this by taking each term and subtracting the previous one. Do you see the pattern?

OpenStudy (anonymous):

no ^

OpenStudy (anonymous):

The rule is +7

OpenStudy (anonymous):

Take each value and subtract the previous one: 14-7=? 21-14=? 28-14=? 35-28=?